# Decision-oriented Text Evaluation

### v1

@@ Abstract
Yu-Shiang Huang1, 2, Chuan-Ju Wang1, Chung-Chi Chen3, 1Academia Sinica, 2National Taiwan University, 3National Institute of Advanced Industrial Science and Technology, Correspondence: c.c.chen@acm.org
@@ Appendix A Comparison of Using Different LLMs for Generating Market Digest
In the main text, we primarily focused our analyses and discussions on texts generated by GPT-4o. To justify this decision, we conducted additional experiments using another large language model, Gemini-2.0-Flash, across an extended dataset covering 89 days (approximately four months‘ trading days). Specifically, we generated market digests (morning briefs and closing-bell reports) using both Gemini-2.0-Flash and GPT-4o, then evaluated these digests using identical large language models (Claude-3-5-Sonnet, Gemini-2.0-Flash, and GPT-4o) as autonomous investors.
Table 3 summarizes these comparative results. The findings consistently indicate superior decision-making efficacy when investors utilize market digests generated by GPT-4o, across both morning briefs and closing-bell report scenarios. Notably, GPT-4o-generated texts led to improved decision accuracy compared to texts generated by Gemini-2.0-Flash, particularly evident in scenarios involving professional-insight-driven asset selection.
Given these clear results demonstrating the effectiveness of GPT-4o-generated texts, our primary human evaluation analyses focused specifically on comparing human investor performance against GPT-4o-generated market digests.

### v2

@@ Abstract
Yu-Shiang Huang,1, 2 Chuan-Ju Wang,1 Chung-Chi Chen3 1Academia Sinica, Taiwan 2National Taiwan University, Taiwan 3National Institute of Advanced Industrial Science and Technology, Japan F09946004@ntu.edu.tw, cjwang@citi.sinica.edu.tw, c.c.chen@acm.org
@@ Appendix A Comparison of Using Different LLMs for Generating Market Digest
In the main text, we primarily focused our analyses and discussions on texts generated by GPT-4o. To justify this decision, we conducted additional experiments using another large language model, Gemini-2.0-Flash, across an extended dataset covering 89 days (approximately four months‘ trading days). Specifically, we generated market digests (morning briefs and closing-bell reports) using both Gemini-2.0-Flash and GPT-4o, then evaluated these digests using identical large language models (Claude-3-5-Sonnet, Gemini-2.0-Flash, and GPT-4o) as autonomous investors. Table 3 summarizes these comparative results.